1. Prove that h(t) * x(t) = x(t) * h(t)
Answer #1

h(t) *XCH) = (hlt-sards, Put t-s= 4 8 = w.. مه حلی ( - 8 = 8 به = ده hits #x15) - - l how)) t-w) dw Grect-w). how dw = x(H *